OpenStudy (anonymous):

What is the domain and range of this function? f(x) = y = log 1/5 x

OpenStudy (anonymous):

Is it: a) \(\large y = log(\frac{1}{5} x)\) b) \(\large y = log(\frac{1}{5x})\) c) \(\large y = log(\frac{1}{5})x\) d) none of the above

OpenStudy (anonymous):

The question is analyze the graph of f(x) = y = log 1/5 x. I typed the function into my calculatr as lg x divided by log 1/5. I have two points that intercept the axis (1,0) and (0,1). I can't find in my notes how to determine the domain and range. I'm sure it is something simple.

OpenStudy (anonymous):

Which function is it though? I'm confused by your notation.

OpenStudy (anonymous):

Domain is all real numbers